### v2

# Less Is More [Uncaptioned image]: Scalable Visual Navigation from Limited Data

Abstract
Imitation learning provides a powerful framework for goal-conditioned visual navigation in mobile robots, enabling obstacle avoidance while respecting human preferences and social norms. However, its effectiveness depends critically on the quality and diversity of training data. In this work, we show how classical geometric planners can be leveraged to generate synthetic trajectories that complement costly human demonstrations. We train Less is More (LiMo), a transformer-based visual navigation policy that predicts goal-conditioned SE(2) trajectories from a single RGB observation, and find that augmenting limited expert demonstrations with planner-generated supervision yields substantial performance gains. Through ablations and complementary qualitative and quantitative analyses, we characterize how dataset scale and diversity affect planning performance. We demonstrate real-robot deployment and argue that robust visual navigation is enabled not by simply collecting more demonstrations, but by strategically curating diverse, high-quality datasets. Our results suggest that scalable, embodiment-specific geometric supervision is a practical path toward data-efficient visual navigation.
Project page: https://leggedrobotics.github.io/less-is-more
I Introduction
Figure 1: Starting from GrandTour [9], a high-quality dataset of recorded quadruped robot missions, we augment the limited real-world expert demonstrations with planner-generated trajectories to scale supervision for imitation learning. The augmented dataset is then used to train LiMo, an end-to-end visual navigation policy.
Navigation on mobile robots has traditionally followed a geometric pipeline: build a map, estimate traversability, and plan a path to the goal. While principled and precise for collision avoidance, this approach struggles when semantic understanding or nuanced behaviors are required. Purely geometric reasoning treats bushes or tall grass as rigid obstacles, yet it can misjudge glare, puddles, or loose gravel as safe. Encoding social norms and complex preferences is likewise difficult in unstructured, real-world settings.
Current work focuses on learning navigation behaviors directly from RGB images using imitation learning, with an emphasis on scaling up both model capacity and dataset size by training on data from different embodiments. The resulting class of models, termed foundation navigation models, has shown promising preliminary results, demonstrating multi-embodiment control and goal-conditioned navigation via poses, images, or language.
Scaling-first approaches have proven particularly effective in autonomous driving, where virtually unlimited human demonstration data can be collected. However, in many real-world robotic settings, data is scarce, necessitating the incorporation of more first-principled approaches into the training of navigation policies. In such settings, transparent and rigorous analysis of model limitations becomes essential for guiding future development.
Under this lens, we identify five core limitations that can pose significant challenges for current visual navigation models: (i) The navigation capability relies on costly real-world expert demonstrations. (ii) Action diversity in existing datasets is limited, biasing navigation toward simple, conservative motion. (iii) Models must learn geometric obstacle avoidance entirely from demonstrations. (iv) Current models trained on cross-embodiment data cannot choose actions that are desirable for a specific embodiment. (v) Training often focuses only on feasible goals.
In response, we ask: How can we build upon the promising imitation learning framework while efficiently supervising navigation policies at scale to acquire both semantic understanding and robust navigation behavior, conditioned on the embodiment?
We propose LiMo, an embodied visual navigation policy that, given a single RGB image and an $\mathrm{SE}(2)$ goal, predicts a sequence of waypoints for local goal-conditioned navigation. To enable scalable supervision, we introduce two complementary datasets derived from GrandTour [9]. The teleop dataset transfers semantic preferences from expert trajectories, while the geometric dataset scales supervision by sampling goals and labeling them with planner-generated trajectories. Because sampled goals may be infeasible, LiMo also learns safe exploratory behavior and robustness to unreachable targets, making it a practical drop-in vision-based local planner.
Across held-out GrandTour missions, LiMo trained on the augmented dataset (teleop + geometric) achieves substantially higher Success weighted by normalized inverse Path Length (SPL) [3] than when trained solely on teleop-data, while preserving strong performance on teleop-style splits. We show that the policy adapts paths to visible structures and selects semantically reasonable shortcuts when safe, while transparently exposing the approach’s current limitations. We further demonstrate real-robot deployment, illustrating LiMo’s practicality as a drop-in replacement for local path planning. We release all datasets and pretrained models to enable reproducibility and further research.
Contributions.
  • LiMo: end-to-end, embodied, goal-conditioned navigation from a single RGB image
  • A scalable imitation learning pipeline combining expert demonstrations and planner-generated supervision
  • Analysis linking path-planning performance to action diversity
  • Closed-loop real-robot deployments of LiMo as a drop-in local planner
II Related Work
II-A Geometric Path Planning
Existing geometric planners either use raw sensory data directly or rely on aggregated spatiotemporal geometric information, such as elevation maps, to compute feasible and safe paths. These planners typically frame path planning as a search or optimization problem. Prominent approaches include graph-based search algorithms [11] and optimization methods minimizing a cost function, such as MPPI [41] or iCEM [30].
These methods are interpretable and allow easy behavior tuning by adjusting the cost function, e.g., to increase clearance to obstacles or change velocity profiles. This, together with formal guarantees of convergence to an optimal path (if it exists), makes them a favored option in industry. However, purely geometric sensing can miss semantically salient cues in unstructured terrain; e.g., unstable materials may appear traversable, while light grass is interpreted as a hard obstacle, missing nuanced navigation behavior. This motivates the use of vision to incorporate semantic understanding into robot navigation [5, 8].
II-B Visual Navigation
Visual navigation aims to learn navigation behaviors directly from camera input. Previous works leverage reinforcement learning (RL) [45] and imitation learning (IL) [23] to train visuomotor control policies.
IL has produced remarkable advances in autonomous driving (AD)  [20, 21, 26]. Lately, several large-scale, high-quality AD datasets have been made available [43, 38]. In contrast, the use of IL for navigating other types of mobile robots is hindered by the limited availability of demonstration data. Collecting expert demonstrations is costly, and existing datasets typically cover only a few hours of operation, which limits robustness and generalization.
Recent goal-image–conditioned models, such as ViNT [34], use transformer-based architectures to learn navigation policies intended to generalize across varied settings and robot embodiments. NoMaD [36] extends ViNT by incorporating a diffusion head for trajectory prediction, which supports goal-conditioned navigation and unconditioned exploration. FlowNav [10] is based on NoMaD but replaces diffusion with conditional flow matching to improve inference efficiency. Additionally, FlowNav uses an off-the-shelf monocular depth estimation model to generate depth priors, which are then fed to the model. These models use a goal image (the observation when the robot is at the goal) for goal-conditioning. This goal image is available during training on pre-recorded robot demonstrations, but it may not be available during deployment.
To overcome the limited data problem, [34, 36, 10] train on a mix of datasets; however, the overall scale remains limited. The total hours reported for ViNT/NoMaD training are $80\text{\,}\mathrm{h}$ across the collected datasets (see Tab. I) and even less for FlowNav, which is modest given the complexity of the task. These works compellingly demonstrate the promise of imitation learning for navigation; nevertheless, the scarcity of expert data continues to hinder generalization.
Also notable is the significant diversity of embodiments in the training data, ranging from slow differential-drive robots operating at $0.5\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ to fast all-terrain vehicles reaching $20\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$. To handle this variation, the models are provided with the past five camera frames to implicitly capture embodiment, and they rely on a normalized action space where predicted waypoints are scaled by each robot’s maximum speed.
However, it is questionable whether cross-embodiment training genuinely improves model performance or whether it is primarily a compromise to increase dataset size. In many cases, training navigation models on embodiment-specific data may be preferable, enabling the full use of a given platform’s motion capabilities. For instance, when training on wheeled-robot data, a quadruped robot might learn overly conservative behaviors, such as avoiding stairs, thereby unnecessarily limiting its mobility and failing to leverage its full potential.
TABLE I: Overview of datasets for imitation learning
Dataset Platform Total [h] Env. Sup.
Expert (ViNT & NoMaD)
BDD-100K [43] Cars $10$ R H
Berkeley [33] Jackal 44 S H
CoryHall [17] RC car 22 H H
GoStanford [16] TurtleBot2 $17$ O H
NeBula [1] ATV $10$ F H
RECON [32] Jackal $25$ F H
SACSoN [15] TurtleBot2 $75$ O H
SCAND-J [18] Jackal 11 W H
SCAND-S [18] Spot 88 W H
Seattle [31] Warthog 11 F H
TartanDrive [39] ATV 77 F H
Others
LeLaN [13] Humans+Robots $129$ D O
CityWalker [22] Humans+Cars $2000$ C O
FrodoBots-2k [14] Earth Rover Zero $2000$ U O
FrodoBots-2k(f) [14] Earth Rover Zero $700$ U O
OmniVLA [12] Cars+Humans+Robots $9500$ D O
TartanGround [28] Robots $40$ D P
Ours
GrandTour [9] ANYmal D 66 D H
Ours ANYmal D $2238$ D P
Env. legend: R=on-road,  S=suburban,  H=hallways,  O=office,  F=off-road,  W=sidewalks,  C=city walking/driving,  U=urban sidewalks/ped. zones,  D=highly diverse (in/out).
Sup.: H=Human Expert, P=Geometric Planner, O=Other.
LiReN [37] approaches data scarcity by continuously improving the model during deployment. They first use offline RL to train a navigation policy on the same multi-embodiment dataset as ViNT. Then they deploy the model and use an actor-critic algorithm to continuously improve it during deployment without needing a human expert to collect more data.
LeLaN [13] uses NoMaD and a vision–language model (VLM) to automatically label first-person view (FPV) YouTube videos for navigation policy learning.
CityWalker [22] also uses a dataset of FPV videos from YouTube, on which it runs visual odometry to automatically generate supervision. The FrodoBots-2k dataset [14] uses crowdsourcing; volunteers remotely drive inexpensive robots (e.g., Earth Rover Zero) to collect demonstrations in urban pedestrian zones. This strategy scales data collection without on-site experts; however, the dataset suffers from poor tracking, inaccurate localization, and low-quality sensors. This is addressed by using Model-Based ReAnnotation (MBRA) [14], an approach that transforms a low-quality dataset into a filtered version with high-quality supervision.
NaVILA [6] scales instruction-following navigation using human egocentric touring videos from YouTube. They estimate camera motion using an off-the-shelf model, generate language descriptions of the tours using a VLM, and train a two-stage vision–language navigation model on the resulting data.
OmniVLA [12] is an omni-modal vision-language-action (VLA) model for navigation that supports goal conditioning via 2D goal poses, egocentric goal images, natural language instructions, and their combinations. Built on top of OpenVLA [19], OmniVLA leverages a large pretrained VLA backbone and is further trained on approximately $9500\text{\,}\mathrm{h}$ of navigation data. This dataset is the combination of [34, 13, 14, 42] datasets from 10 different robotic platforms. Notably, the dataset mixture is heavily skewed towards BDD-V [42], i.e., on-road autonomous driving data. They use a customized MBRA-style [14] approach to generate synthetic, feasible supervision for their target robots on BDD-V.
Instead of predicting waypoints directly, VENTURA [44] uses an internet-pretrained image diffusion model to generate an image-space path mask conditioned on the current RGB observation and a language instruction. Another model then converts the image-based path mask into waypoints. To scale supervision without manual annotations or accurate odometry, VENTURA generates path-mask labels indicating where the robot moved during deployment from egocentric videos using point tracking and adds captions using a VLM.
While [13, 6, 12, 44] study language-conditioned navigation, we focus only on goal-conditioned visual navigation. Despite strong progress in visuomotor policies, geometric reasoning in cluttered environments remains challenging, likely reflecting the limited geometric diversity of current datasets, which are often dominated by AD or conservative motion.
LoGoPlanner [29] addresses this by grounding navigation in metric-scale geometry through auxiliary localization and reconstruction tasks. In contrast, we scale embodiment-specific geometric supervision by augmenting a high-quality legged-robot dataset with planner-generated trajectories, achieving geometry-aware performance without explicit 3D reconstruction at deployment. Our work extends the GrandTour [9] dataset’s $6\text{\,}\mathrm{h}$ of deployments to thousands of hours of diverse, high-quality, embodiment-specific trajectories.
III Method
Figure 2: Three representative samples from the augmented dataset $\mathcal{D}_{\text{AUG}}$. Each row shows the front camera image (left) and the corresponding elevation map (right). Teleoperated, geometric, and real-world paths are overlaid with their associated goals.
III-A Overview
Given an RGB image $\mathbf{I}$ and a goal pose $\mathbf{g}=[x,y,\theta]\in\mathrm{SE}(2)$ in the robot-centric frame, the objective is to predict a path $\mathbf{s}_{1:N}$ of $N$ waypoints $\mathbf{s}_{i}=[x_{i},y_{i},\theta_{i}]\in\mathrm{SE}(2)$ that safely guides the robot towards the goal. We address this by training a policy $\pi:(\mathbf{I},\mathbf{g})\mapsto\mathbf{s}_{1:N}$ via IL, which we call LiMo.
From a base dataset $\mathcal{D}_{\text{Base}}$ containing time-synchronized RGB video and elevation maps from expert teleoperated deployments, we extract front-camera images paired with executed $\mathrm{SE}(2)$ path segments to form the teleop dataset $\mathcal{D}_{\text{TEL}}$ (Sec. III-C1). We then sample additional goals per frame and label them with trajectories planned using a geometric MPPI planner (Secs. III-B and III-C2), yielding the geometric dataset $\mathcal{D}_{\text{GEO}}$. Finally, we train LiMo on the augmented dataset $\mathcal{D}_{\text{AUG}}=\mathcal{D}_{\text{TEL}}\cup\mathcal{D}_{\text{GEO}}$ (Sec. III-D).
III-B MPPI Planning
MPPI is a gradient-free optimization algorithm used for robot motion and path planning. Originally introduced in [41], MPPI’s effectiveness has been demonstrated in several real-world experiments, such as high-speed driving and whole-body control of legged robots [41, 2]. We use MPPI to solve the following optimization problem:
$\min_{\mathbf{a}_{0:N-1}}\quad J(\mathbf{s}_{0},\mathbf{a}_{0:N-1}),$ (1)
where $\mathbf{a}_{0:N-1}$ is a sequence of command velocities $\mathbf{a}_{i}=[v_{i}^{x},v_{i}^{y},\omega_{i}]$, with $v_{i}^{x},v_{i}^{y},\omega_{i}$ representing the linear and angular velocities, respectively, and $J:(\mathbf{s}_{0},\mathbf{a}_{0:N-1})\mapsto c\in\mathbb{R}$ denoting a predefined cost function that encourages goal-reaching and safety, conditioned on the current state $\mathbf{s}_{0}\in\mathrm{SE}(2)$.
MPPI solves the optimization by iteratively sampling a population of command velocity sequences $\mathbf{a}_{0:N-1}$. It then uses $J$ to weight the population members and update the population distribution, preferring low-cost members.
The future states are computed based on the current state $\mathbf{s}_{0}$ and a dynamics model $\mathbf{s}_{i+1}=f(\mathbf{s}_{i},\mathbf{a}_{i})$. The dynamics model $f$ can be learned or classically derived. In our case, we use a simple piecewise-constant velocity model in $\mathrm{SE}(2)$:
$\theta_{i+1}$ $=\theta_{i}+\omega_{i}\,\Delta t,$ (2)
$\begin{bmatrix}x_{i+1}\\ y_{i+1}\end{bmatrix}$ $=\begin{bmatrix}x_{i}\\ y_{i}\end{bmatrix}+\mathbf{R}(\theta_{i})\begin{bmatrix}v_{i}^{x}\\ v_{i}^{y}\end{bmatrix}\Delta t,$ (3)
where $\mathbf{R}(\theta_{i})\in\mathrm{SO}(2)$ is the planar rotation matrix that maps body-frame velocities to the world frame and $\Delta t$ is the control timestep.
To encourage goal reaching while ensuring robot safety, we construct the cost function as follows:
$\begin{split}J(\mathbf{s}_{0},\mathbf{a}_{0:N-1})&=w_{\text{trav}}C_{\text{trav}}(\mathbf{s}_{1:N})+w_{\text{goal}}C_{\text{goal}}(\mathbf{s}_{1:N})\\ &\quad+w_{\text{effort}}C_{\text{effort}}(\mathbf{a}_{0:N-1}).\end{split}$ (4)
III-B1 Traversability Cost ($C_{\text{trav}}$)
We use the convolutional neural network (CNN) from [25, 7] to compute a traversability map $\mathbf{T}$ from the elevation map $\mathbf{E}$. The CNN is trained to provide an embodiment-specific traversability estimate based on the robot’s capabilities. The traversability cost is the sum of the grid cells of $\mathbf{T}$ traversed by the robot along $\mathbf{s}_{1:N}$.
III-B2 Goal distance cost ($C_{\text{goal}}$)
By thresholding $\mathbf{T}$, we build a binary obstacle grid map $\mathbf{O}$. From $\mathbf{O}$, a geodesic distance field to the goal is derived [4]. Then $C_{\text{goal}}$ is the sum of the geodesic distances of each waypoint to the goal, encouraging fast goal-reaching. When close to the goal, the heading difference $|\theta_{g}-\theta_{i}|$ is added for goal alignment.
III-B3 Effort cost ($C_{\text{effort}}$)
This is the weighted sum of the linear, lateral, and angular velocities $w_{\text{lin}}|v^{x}_{i}|+w_{\text{lat}}|v^{y}_{i}|+w_{\text{ang}}|\omega_{i}|$ over the command sequence, encouraging efficient motion while discouraging unnecessary sideways motion and excessive turning.
III-C Dataset Curation
III-C1 Teleoperation Dataset ($\mathcal{D}_{\text{TEL}}$)
$\mathcal{D}_{\text{TEL}}$ consists of fixed-horizon, goal-conditioned segments extracted from the recorded teleoperated trajectories and serves to transfer semantic preferences and motion characteristics to the policy.
III-C2 Geometric Dataset ($\mathcal{D}_{\text{GEO}}$)
To increase the diversity and scale of the training data, we employ the previously introduced MPPI-planner to generate additional demonstrations. For every camera frame $\mathbf{I}$, we independently sample $K$ random goals $\mathbf{g}_{1},...,\mathbf{g}_{K}$ from a Gaussian distribution in the robot-centric coordinate frame (see Tab. III in the Appendix). Given the matching elevation map $\mathbf{E}$ and the goals, MPPI plans the paths $\mathbf{s}_{1:N}^{1},...,\mathbf{s}_{1:N}^{K}$; generating $K$ diverse training samples $(\mathbf{I},\mathbf{g}_{k},\mathbf{s}_{1:N}^{k})$ that are saved in $\mathcal{D}_{\text{GEO}}$.
The MPPI planner reliably generates geometrically feasible paths by leveraging privileged information from accurate elevation mapping. This enables the creation of demonstrations that are more diverse and exploratory than those obtained from real-world teleoperation alone. Even when a randomly sampled goal is infeasible, the planner still produces safe, embodiment-consistent behavior, either progressing toward the goal in a best-effort manner or exploring alternative routes while avoiding obstacles (see Fig. 2).
III-D Policy Architecture
Figure 3: Overview of LiMo’s architecture. The policy takes as input a single RGB image $\mathbf{I}$, and a goal pose $\mathbf{g}=(x,y,\theta)$ in the robot-centric frame. Image features are extracted using a frozen DINOv2 [27] encoder and combined with learned positional embeddings. A transformer decoder conditioned on the goal embedding predicts NN waypoint embeddings, which are linearly projected to $N$ waypoints $(x,y,\theta)$, forming the output trajectory.
$\pi$ receives a single RGB image $\mathbf{I}$ together with a $\mathrm{SE}(2)$ goal $\mathbf{g}$ in the robot-centric coordinate frame. It outputs $N$ waypoints in $\mathrm{SE}(2)$, also in the robot-centric frame.
First, a DINOv2 [27] image encoder is used to extract patch embeddings from the image $\mathbf{I}$. The goal pose $\mathbf{g}$ is projected to the same embedding dimension and replicated $N$ times. We call these $N$ goal embeddings waypoint queries.
To every patch embedding and every waypoint query, we add an individually learned absolute positional embedding. Next, the waypoint queries are passed to a 4-block transformer decoder, with the patch embeddings used as keys and values.
The decoder outputs $N$ waypoint embeddings, which are projected to $N$ waypoints via a linear layer with weights shared across all positions. The resulting waypoint sequence forms the predicted plan. An overview of the architecture is shown in Fig. 3.
Further implementation details are available in Sec. -A.
IV Experiments
In this section, we address the following questions: (Q1) Do the diverse geometric demonstrations effectively improve navigation performance? (Q2) Can LiMo generalize to unseen environments? (Q3) Does our method enable embodiment-specific navigation? (Q4) Does LiMo work when deployed on a real robot?
IV-A Datasets
The experiments build on the GrandTour dataset [9], which comprises $6\text{\,}\mathrm{h}$ of highly diverse deployments of an ANYmal D robot across 49 sites in Switzerland, with varying weather and lighting conditions. From these missions, we derive three datasets, in accordance with Secs. III-C1 and III-C2: (i) the teleoperation dataset $\mathcal{D}_{\text{TEL}}$, (ii) the geometric-planner-labeled dataset $\mathcal{D}_{\text{GEO}}$ with $K=10$ trajectories per image, each with $N=50$ waypoints evenly spaced over $T=$ $5\text{\,}\mathrm{s}$, and finally, (iii) the augmented dataset $\mathcal{D}_{\text{AUG}}=\mathcal{D}_{\text{TEL}}\cup\mathcal{D}_{\text{GEO}}$. For evaluation, we reserve six GrandTour missions exclusively as test scenes (see Tab. III in the Appendix for further details on dataset size).
IV-B Qualitative Performance and Emergent Behavior
Figure 4: Qualitative comparison of predicted trajectories across diverse indoor and outdoor scenarios. Images (b, c, e, f, h, i) are included in the training set, while (a, d, g, j, k, m) are held-out scenes. Column three compares predictions from the teleop-only policy with those from the augmented-trained policy, where the latter shows much better scene understanding. Column four illustrates the remaining failure cases. LiMo correctly recognizes door openings (a, j), handles stair ascent and descent consistent with ANYmal’s embodiment (e, i), and navigates unstructured natural terrain such as forests and fields (c, f). It also demonstrates robustness on construction sites (g) and in high-elevation outdoor environments (k), while remaining challenged by cliff edges (d), transparent surfaces, and obstacles such as glass doors and nets (h, m), and by unusual obstacles such as snow piles (m).
In Fig. 4, we present representative path predictions from LiMo. Examples (a, b, e, f, g, i, j, k) highlight a strong geometric understanding and embodiment-aware behavior: LiMo identifies door openings, plans to climb and descend stairs, and navigates rough natural terrain. Comparisons in column three (c, g, k) show that the $\mathcal{D}_{\text{AUG}}$-trained policy produces substantially more geometry-aware and obstacle-avoiding plans than the $\mathcal{D}_{\text{TEL}}$-only baseline, clearly indicating the benefit of including geometric demonstrations (Q1). The model also plans sensible trajectories in held-out missions (a, d, g, j, k), demonstrating generalization to unseen environments (Q2), and consistently chooses behaviors aligned with ANYmal’s locomotion capabilities, confirming embodiment-specific navigation (Q3).
IV-C Quantitative Results
IV-C1 Metrics
Following the recommendations of [3], we evaluate trajectories on held-out missions using geodesic distances (GD) and Success weighted by normalized inverse Path Length (SPL). A path is deemed successful if (i) the GD of the final waypoint to the goal is $\leq$1.0\text{\,}\mathrm{m}$\$ and (ii) the robot does not collide. Collisions and geodesic distances are computed using traversability maps derived from the ground-truth elevation maps. Let $S_{i}\in\{0,1\}$ indicate success for sample $i$, $\ell_{i}$ be the geodesic start-to-goal distance, and $p_{i}$ the path length; then the SPL is defined as:
$\mathrm{SPL}=\frac{1}{M}\sum_{i=1}^{M}S_{i}\,\frac{\ell_{i}}{\max(p_{i},\ell_{i})}\,.$ (5)
IV-C2 Baselines
Geometric Planner Oracle
The path obtained by running the planner described in Sec. III-B. The geometric planner should be viewed as an approximate oracle rather than a deployable baseline. It plans with access to the ground-truth elevation maps that are not available to the learned policies.
Moreover, the planner is used to define the evaluation set: we consider only goals that it can reach, so as not to penalize methods for inherently infeasible targets.
Straight-Line Paths
Paths going straight from the robot to the goals, capped at $5\text{\,}\mathrm{m}$ in length.
Real-World Paths
The real-world paths recorded during dataset collection, only available for $\mathcal{D}_{\text{TEL}}$ goals.
OmniVLA (open-loop) [12]
OmniVLA is a state-of-the-art visual navigation model initialized using the authors’ released checkpoint. Our benchmark measures open-loop, single-shot trajectory prediction for local goal-conditioned planning, whereas OmniVLA is primarily designed for closed-loop replanning. We therefore evaluate OmniVLA under this open-loop protocol, and the results should be interpreted accordingly.
IV-C3 Results
TABLE II: Evaluation summary sorted by SPL
Planner Col. (%) $\Downarrow$ Succ. (%) $\Uparrow$ SPL (%) $\Uparrow$
Evaluated on $\mathcal{D}_{\text{TEL}}$
OmniVLA (open-loop) $12.6$ $67.2$ $41.5$
Trained on $\mathcal{D}_{\text{TEL}}$ $10.8$ $87.1$ $84.4$
Trained on $\mathcal{D}_{\text{AUG}}$ $11.1$ $88.7$ $86.4$
Straight-Line Paths $12.5$ $87.5$ $87.5$
Real-World Paths $11.0$ $89.0$ $87.5$
Geometric Planner $3.7$ $96.3$ $88.8$
Evaluated on $\mathcal{D}_{\text{AUG}}$
OmniVLA (open-loop) $14.2$ $46.5$ $33.5$
Trained on $\mathcal{D}_{\text{TEL}}$ $14.1$ $51.4$ $49.7$
Straight-Line Paths $23.1$ $76.9$ $76.9$
Trained on $\mathcal{D}_{\text{AUG}}$ $14.5$ $82.2$ $80.0$
Geometric Planner $1.0$ $99.0$ $95.4$
As shown in Tab. II, incorporating geometric demonstrations yields substantial SPL improvements on the diverse $\mathcal{D}_{\text{AUG}}$ test set, indicating that geometric supervision significantly strengthens performance in complex, unseen scenarios (Q1, Q2).
Since SPL explicitly rewards efficient, short paths, the Straight-Line baseline performs well on $\mathcal{D}_{\text{TEL}}$, where many goals lie roughly straight ahead, and few obstacles are encountered. This highlights that geometric augmentation is most beneficial in diverse, obstacle-rich scenarios, where simple straight-line behavior no longer suffices.
The non-zero collision rate of the Real-World Paths baseline mainly reflects limitations of the traversability map based collision proxy. Collisions are determined from elevation maps, which can label rough but physically traversable terrain as obstacles. As a result, safely executed recordings are sometimes counted as collisions. This makes the reported collision rates conservative and biased toward overestimating failures.
We observe that OmniVLA achieves competitive collision avoidance, even on the more challenging $\mathcal{D}_{\text{AUG}}$ test set, but exhibits reduced goal-reaching performance. This is expected in part due to the evaluation protocol: OmniVLA is trained for closed-loop replanning with fixed-length waypoint chunks rather than open-loop trajectory termination at a specified goal.
In addition, we hypothesize that training on large, cross-embodiment datasets, dominated by wheeled platforms, can further limit goal-reaching accuracy for a specific embodiment, as the model must implicitly infer platform dynamics and feasible motions solely from visual context. Finally, the limited geometric diversity of legged-robot data within such datasets may restrict the model’s exposure to embodiment-specific maneuvers (e.g., lateral or in-place motions). A more detailed discussion of these factors is provided in Sec. -B.
IV-D Ablations
Figure 5: (Left) Evaluation of SPL on $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{AUG}}$ test sets trained on a varying number of augmenting geometric paths. Adding geometric samples mainly benefits performance on the more diverse $\mathcal{D}_{\text{AUG}}$ test split. (Middle) When evaluating goals following the $\mathcal{D}_{\text{AUG}}$ distribution, policies trained solely on $\mathcal{D}_{\text{TEL}}$ consistently underperform compared to those trained on the $\mathcal{D}_{\text{AUG}}$. While training on $\mathcal{D}_{\text{AUG}}$ generalizes well to $\mathcal{D}_{\text{TEL}}$-style goals, policies trained only on $\mathcal{D}_{\text{TEL}}$ fail to generalize to unseen diverse goals with limited data. (Right) Evaluation of SPL on $\mathcal{D}_{\text{TEL}}$ goals trained using the $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{AUG}}$. Training on just 5 missions of $\mathcal{D}_{\text{AUG}}$ achieves comparable performance to training on $\mathcal{D}_{\text{TEL}}$ with 41 missions.
IV-D1 Training with fewer Geometric Paths per Frame
To explore the benefits of adding geometric paths, we train policies on different mixtures of $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{GEO}}$. The policies are evaluated on the held-out test missions with $\mathcal{D}_{\text{TEL}}$ or more diverse $\mathcal{D}_{\text{AUG}}$ distributed goals (see Fig. 5 (left)).
IV-D2 Training on fewer Missions
To understand how well policies generalize when trained with fewer missions, we compare performance across varying numbers of training missions drawn from either $\mathcal{D}_{\text{TEL}}$ or $\mathcal{D}_{\text{AUG}}$. Policies trained on a small number of diverse $\mathcal{D}_{\text{AUG}}$ missions retain strong performance on $\mathcal{D}_{\text{TEL}}$ goals, whereas policies trained only on $\mathcal{D}_{\text{TEL}}$ require substantially more missions to reach comparable SPL (see Fig. 5). This highlights that diversity in the training data, rather than sheer quantity, is key for efficient generalization (Q1, Q2). Tab. III in the Appendix illustrates the difference in goal diversity among datasets.
IV-E Robot Deployment
We deploy LiMo on an ANYmal D quadruped robot in locations outside the training scenes to demonstrate its usability in real-world closed-loop control (Q4), embodiment-specific behavior (Q3), and robustness in unseen, cluttered environments (Q2). These experiments illustrate the robot clearing obstacles, following corridors, and reacting to dynamic scene changes.
We omit on-robot comparisons to prior end-to-end visual navigation models. Existing approaches are designed for cross-embodiment deployment and often do not support pose-based goal conditioning. In contrast, LiMo predicts high-resolution, goal-pose-conditioned $\mathrm{SE}(2)$ trajectories that terminate at the commanded goal and are tailored to ANYmal’s embodiment.
Moreover, representative models such as OmniVLA [12] are substantially larger, cannot be executed at comparable rates on the robot’s onboard hardware, and rely on additional mechanisms such as topological maps to provide subgoals. We therefore deliberately choose to only highlight the real-world performance of our most powerful model.
IV-E1 Deployment Setup
We run LiMo on a NVIDIA Jetson Orin on-board the robot in closed loop at $6\text{\,}\mathrm{H}\mathrm{z}$. We use a simple lookahead path follower node to make the robot track the predicted $\mathrm{SE}(2)$ waypoints.
IV-E2 Incorporating Side Cameras
Using only the front-facing camera in closed-loop deployment led to collisions, as obstacles were no longer detected once they left the forward field of view. Therefore, in the robot experiments, we use a multi-camera variant with additional left- and right-facing cameras, implemented via a minimal architectural extension.
More details are provided in Sec. -C.
IV-F Obstacle Course Navigation
Figure 6: LiMo navigating an indoor obstacle course shown at four different time steps. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory is shown in cyan and the goal is shown in green. LiMo generates smooth, collision-free paths while adapting to obstacles. The LiDAR data is recorded for visualization purposes only and is not used as input to the policy.
We design an indoor obstacle course comprising rigid obstacles and small traversable elements, requiring the robot to plan non-trivial paths while exploiting its legged mobility. The robot is given a single local goal and must navigate around obstacles without collisions (see Fig. 6). The experiment demonstrates that LiMo produces smooth, geometry-aware paths. Despite operating purely on vision at inference time, the predicted trajectories remain consistent with the robot’s footprint and motion capabilities.
IV-G Reacting to Dynamic Obstacles
Figure 7: LiMo reacting to a dynamic obstacle (a person carrying a red mat). Rows show successive time steps during execution. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory (cyan) adapts online as the obstacle blocks the original path, leading to safe avoidance behavior.
To test reactivity, we introduce a moving obstacle by having a person carry a large mat and intentionally block the robot’s path during execution. LiMo runs continuously at $6\text{\,}\mathrm{H}\mathrm{z}$, allowing the robot to replan as the obstacle moves.
As shown in Fig. 7, the robot quickly deviates from its original path and selects alternative routes, including lateral and backward embodiment-specific motion, to avoid colliding. This behavior demonstrates that LiMo can handle dynamic environmental changes.
IV-H Corridor Following
We deploy the robot in a long corridor scenario with limited lateral clearance. The robot must stay centered as it progresses toward a goal located to the right of the corridor exit (see Fig. 10 in the Appendix).
V Limitations
Planner supervision depends on elevation and traversability estimates and can fail in edge cases, leading to unsafe or suboptimal labels. At inference, LiMo operates on a single RGB frame without temporal memory, which limits reasoning under occlusions and causes obstacles to be forgotten once they leave the field of view. While geometric augmentation scales supervision efficiently, visual diversity is still constrained by the underlying dataset. Fig. 4 illustrates representative failure cases, and a continuing discussion is provided in Sec. -D.
VI Conclusion
We presented LiMo, a goal-conditioned visual navigation policy trained on a mix of teleoperation data and scalable geometric supervision. By combining expert demonstrations with planner-generated trajectories that are fully embodiment-specific, we scale supervision to thousands of hours of high-quality trajectories. Across held-out missions, geometric augmentation leads to clear improvements in SPL, especially in diverse, obstacle-rich environments. LiMo learns geometry-aware, semantically grounded behavior aligned with ANYmal’s embodiment, and experiments on a real robot validate its utility as a practical path planner.
Overall, our results show that for vision-based goal-conditioned navigation, curating diverse, geometry-aware supervision can be more effective than relying solely on large, heterogeneous datasets.
By combining limited expert demonstrations with scalable, embodiment-specific geometric supervision, we provide a practical and data-efficient path toward robust visual navigation in real-world robotic systems.
Acknowledgments
The authors would like to thank Turcan Tuna and William Talbot for their support during robot experiments and system integration. We also thank members of the Robotic Systems Lab at ETH Zurich for valuable discussions and feedback throughout the project. We thank the authors of OmniVLA for providing model checkpoints, code snippets, and helpful clarifications.
Fig. 8. Visualization of goal distributions based on 10,000 samples from each dataset. (Left) Goals from the teleoperation dataset $\mathcal{D}_{\text{TEL}}$ (blue), sampled along the recorded trajectories. (Right) Goals from the augmented dataset $\mathcal{D}_{\text{AUG}}$, including teleoperation goals (blue) and additional randomly sampled goals (orange).
TABLE III: Dataset statistics
Dataset Split #Samples Length [m] Time [h] Avg. vel. [m/s]
$\mathcal{D}_{\text{TEL}}$ Train 113 326113\,326 281 870281\,870 $157$ $0.50$
Test 17 13917\,139 43 21243\,212 $24$ $0.50$
$\mathcal{D}_{\text{GEO}}$ Train 1 289 9301\,289\,930 4 457 7294\,457\,729 $1792$ $0.69$
Test 191 060191\,060 653 978653\,978 $265$ $0.68$
$\mathcal{D}_{\text{AUG}}$ Train 1 403 2561\,403\,256 4 739 5994\,739\,599 $1949$ $0.68$
Test 208 199208\,199 697 190697\,190 $289$ $0.67$
Each sample is a fixed-horizon waypoint sequence with duration $T=$5\text{\,}\mathrm{s}$\$. The reported path length is the sum of Euclidean distances between consecutive waypoints. The reported average velocity should be interpreted with care: while the robot typically walks at around $1\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$, many samples include stopping behavior (e.g., trajectories that reach the goal early), which lowers the mean speed.
Figure 9: Comparison of trajectory predictions with and without side-facing cameras. Five goals (green) and the corresponding predicted trajectories are shown. Using only the front camera (cyan), LiMo loses awareness of a partially cleared obstacle once it leaves the forward field of view, leading to collisions. With left and right side camera views (red), obstacle awareness is preserved, and all paths remain collision-free.
Figure 10: LiMo is following a narrow corridor toward a goal located at the far end. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory (cyan) remains centered and collision-free despite tight lateral constraints.